The line through (8,-2) perpendicular to the line y=7-2x

What an equation of the line described

Answer by lwsshak3(11628)

The line through (8,-2) perpendicular to the line y=7-2x
Write an equation of the line described
**
Equation of a line: y=mx+b, m=slope, b=y-intercept
y=7-2x
y=-2x+7
m=-2
slope of the perpendicular line=1/2 (negative reciprocal)
y=x/2+b
solve for b using coordinates of given point on the line (8,-2)
-2=8/2+b
b=-6
Equation of the perpendicular line:y=x/2-6